Chapter 13 (13.104)

Carbon disulfide (CS2) boils at 46.30 ∘C and has a density of 1.261 g/mL.

1) When 0.250 mol of a nondissociating solute is dissolved in 390.0 mL of CS2, the solution boils at 47.49 ∘C. What is the molal boiling-point-elevation constant for CS2?

2) When 5.39 g of a nondissociating unknown is dissolved in 50.0 mL of CS2, the solution boils at 47.08 ∘C. What is the molecular weight of the unknown?
1
Expert's answer
2016-11-11T10:39:15-0500
1) The difference in boiling point is proportional to the concentration of a dissolved compound. Thus, E = (t(sol) - t(CS2))*V(CS2)*d(CS2)/n(sol) = 2341 C*g/mol.
2) Now, we can use this constant to calculate the molar mass of the unknown: M = E*m(unk.)/(t(unk.) - t(CS2))/m(CS2) = 257 g/mol.